# SciNLP: A Domain-Specific Benchmark for Full-Text Scientific Entity and Relation Extraction in NLP

### v3

@@ Abstract
Structured information extraction from scientific literature is crucial for capturing core concepts and emerging trends in specialized fields. While existing datasets aid model development, most focus on specific publication sections due to domain complexity and the high cost of annotating scientific texts. To address this limitation, we introduce SciNLP—a specialized benchmark for full-text entity and relation extraction in the Natural Language Processing (NLP) domain. The dataset comprises 60 manually annotated full-text NLP publications, covering 7,072 entities and 1,826 relations. Compared to existing research, SciNLP is the first dataset providing full-text annotations of entities and their relationships in the NLP domain. To validate the effectiveness of SciNLP, we conducted comparative experiments with similar datasets and evaluated the performance of state-of-the-art supervised models on this dataset. Results reveal varying extraction capabilities of existing models across academic texts of different lengths. Cross-comparisons with existing datasets show that SciNLP achieves significant performance improvements on certain baseline models. Using models trained on SciNLP, we implemented automatic construction of a fine-grained knowledge graph for the NLP domain. Our KG has an average node degree of 3.2 per entity, indicating rich semantic topological information that enhances downstream applications. The dataset is publicly available at: https://github.com/AKADDC/SciNLP.
@@ 1 Introduction
Scientific Named Entity Recognition (SciNER) and Scientific Relation Extraction (SciRE) have emerged as two pivotal technologies for scientific literature mining. SciNER focuses on identifying and categorizing domain-specific scientific entities in publications(Jeong and Kim, 2022; Zhang et al., 2021; Zhong and Chen, 2021a; Beltagy et al., 2019), while SciRE detects semantic relationships between entities and classifies them into predefined relation types(Yan et al., 2021; Chen et al., 2020; Ji et al., 2020; Wadden et al., 2019). These tasks play essential roles in downstream applications such as scientific leaderboard construction, academic Q&A systems, knowledge graph building, text summarization, and method recommendation(Wang et al., 2024; Cao et al., 2021; Zhang et al., 2024a; Kedia et al., 2021).
@@
We construct SciNLP, the first manually curated, full-text scientific publication dataset for the NLP domain, comprising 60 ACL conference long papers (2001–2024) with comprehensive semantic annotations. The dataset includes 7,072 fine-grained entities and 1,826 complex relations, addressing the limitations of existing scientific information extraction resources confined to abstracts or narrow subdomains. SciNLP pioneers systematic full-text entity annotation for NLP literature.
@@ 2.2 Scientific Entity Relation Extraction
Reviewing the extant annotated datasets reveals significant imbalances in disciplinary coverage and annotation depth across current scientific publication resources for SciIE. Cross-domain corpora such as predominantly concentrate on AI and ML domains(Luan et al., 2018; Pan et al., 2024; Zhang et al., 2024b), while specialized disciplinary datasets exhibit marked fragmentation in textual coverage, typically constrained to specific paper sections or isolated sentences(Jain et al., 2020; Gábor et al., 2018), thereby compromising the integrity of holistic scientific discourse. Addressing these limitations, this study proposes the creation of SciNLP – the inaugural full-text annotation benchmark for NLP research. Anchored in comprehensive semantic units from 60 ACL proceedings (2001-2024), the corpus achieves granular annotation of 7,072 entities and 1,826 relations. Furthermore, it establishes a sophisticated taxonomy of entity and relational categories specifically tailored to the structural and conceptual composition of NLP scholarly articles.
@@ 3.2 Comparison with Previous Datasets
Table 2: Examples of the relations types, the four entity types: Method, Metric, Dataset, Task.
@@ 3.3 Data Analysis
Upon completing the annotation process, our dataset comprises 7,072 fine-grained entities and 1,826 relations, with an average of 121 entities and 31 relations per document. As shown in Table 1, the scale of our dataset substantially exceeds prior NLP-specific datasets designed for entity and relation extraction. The annotated data was then randomly shuffled and split into training, validation, and testing sets in an 8:1:1 ratio to support subsequent experimental analyses.
@@ 4.2 Evaluation Settings and Results
Boundary-Aware Metric (Rel): Following prior work(Zhong and Chen, 2021b; Ye et al., 2022; Yan et al., 2023), a relation is considered correct only if: (a) Subject entity boundaries are correctly identified; (b) Object entity boundaries are correctly identified; (c) Relation type is accurately predicted.
@@
PURE(Zhong and Chen, 2021b) 59.42 57.51 56.15 57.03 60.25 59.93 58.35 59.93
@@
HGERE 94.15 61.74 60.76 - 74.91 55.72 43.6 - 86.85 62.32 61.10 -
Table 4: F1 scores of different baselines on our proposed dataset.
@@ 5.1 Supervised Baselines
PURE: (Zhong and Chen, 2021b) This model employs a dual-encoder architecture to decouple entity recognition and relation extraction tasks. The entity model performs span-based type prediction, while the relation model enhances classification through dynamic entity pair inputs and the insertion of entity-type markers. It introduces crosssentence context expansion by extending input windows, leveraging pre-trained models’ ability to capture long-range dependencies for improved global semantic reasoning.
@@ 5.2 Result Analysis
Table  4 shows a comparison between SciNLP and scientific datasets like SCIERC and SciER, and it reveals critical differences and their impact on model performance: First, SciNLP’s significantly smaller entity annotation count compared to SciER results in a notable performance gap for traditional sequence labeling models like PURE in NER tasks (SciNLP vs. SciER). However, SciNLP achieves comparable NER performance to SCIERC, validating the strong dependency of sequence models on training data volume and their performance ceilings under annotation scale constraints. Furthermore, SciNLP and SciER exhibit distinct advantages in complex relation modeling for the Rel+ task. For instance, HGERE underperforms on SCIERC’s Rel+ task compared to SciNLP and SciER, a disparity attributable to SciNLP/SciER’s cross-document triplet annotation framework, which deeply encodes scientific semantic networks through interdisciplinary knowledge linkages. These findings highlight how annotation design, particularly full-text and domain-specific strategies, fundamentally shapes model capabilities in capturing scientific knowledge structures.
Cross-domain transfer experiments further reveal SciNLP’s domain-specific specialization. When transferring the HGERE model from SciNLP to SCIERC, its NER performance degrades significantly—a phenomenon attributable to systematic differences in disciplinary focus: SciNLP employs NLP-centric deep annotation, while SCIERC adopts a cross-disciplinary paradigm. Experiments confirm that domain-focused annotation strategies effectively enhance intra-domain precision—under identical model architectures, SciNLP’s disciplinary specialization achieves 92.4 entity recognition precision in computational linguistics, far surpassing cross-disciplinary scenarios. This "depth vs. breadth" trade-off demonstrates that specialized annotation frameworks reduce inter-domain distribution shifts, significantly improving models’ ability to capture domain-specific semantic patterns, albeit at the cost of cross-domain generalization capabilities.
@@ 6 Building Fine-grained KG in NLP Domain
We construct a scientific knowledge graph from the entire ACL corpus. First, we built a comprehensive corpus containing full texts and metadata of all ACL conference papers (2001–2024) from the ACL Anthology repository, totaling 82,672 papers. Leveraging the HGERE model, we performed end-to-end triplet extraction across all full-text papers to build a fine-grained NLP knowledge graph, where nodes represent extracted scientific entities and edges denote semantic relationships between entities. This framework enables systematic encoding of domain-specific knowledge structures and cross-document scientific logic.
The construction of our knowledge graph involves two key steps: First, the model extracts all triplet data T from each paper, preserving every triplet in the form ($e_{i}$, $r_{k}$, $e_{j}$), where $e_{i}$ and $e_{j}$ are related entities, and $r_{k}$ denotes their semantic relationship. Second, we perform entity normalization to ensure uniqueness within the knowledge graph. This process begins by creating an abbreviation-full form mapping table between entities using the predefined similar relationship. Subsequently, we unify entities using the entity normalization method proposed by(Zhang et al., 2024a), which calculates similarity scores based on lexical features and applies clustering algorithms to standardize entity representations. This dual-step approach guarantees both the richness and semantic consistency of the knowledge graph. In the end, we constructed a fine-grained NLP knowledge graph comprising 232k nodes and 743k edges, with an average node degree of 3.2. Figure  2 illustrates a subset of the constructed graph, demonstrating its ability to effectively preserve domain-specific knowledge in NLP. This topology reflects the intricate semantic networks inherent to computational linguistics research.
@@
Task 38,129 measuredBy 125,494
Dataset 35,756 evaluatedBy 26,080
Metric 34,809 usedFor 184,334
Model 124,106 enhancedBy 34660
evaluatedOn 40,465
partOf 47,669
compareWith 82,599
subtaskOf 43,372
similarWith 40,689
trainedWith 97,723
subclassOf 20,179
Total 232,800 Total 743,264
@@ 7 Conclusion
We present SciNLP, a dataset specifically designed for fine-grained entity and relation extraction in the Natural Language Processing (NLP) domain. The dataset defines four entity types (Task, Method, Dataset, Metric) and 11 relation types (MeasuredBy, UsedFor, TrainedWith, EnhancedBy, SubclassOf, SubtaskOf, PartOf, EvaluatedBy, EvaluatedOn, SimilarWith, CompareWith), overcoming the limitations of coarse-grained relation definitions in existing resources. Constructed from 60 ACL conference papers spanning 2001–2024, SciNLP contains 7,072 entities and 1,826 relations. Unlike cross-disciplinary or abstract-level datasets, SciNLP preserves full-text scientific context through document-level annotation, providing critical support for models to capture cross-paragraph semantic dependencies. Experimental results demonstrate that models trained on SciNLP significantly outperform sentence-level settings in document-context modeling. For instance, the HGERE model achieves a 17.16 point improvement on the strict relation classification (Rel+) task compared to the cross-disciplinary SciERC dataset, validating the enhanced capability of domain-specific annotation for complex relation reasoning. Leveraging SciNLP, we automatically extracted knowledge from 82,672 NLP publications, constructing a fine-grained knowledge graph with 232k nodes and 743k edges, where entities exhibit an average node degree of 3.2. SciNLP establishes a high-confidence benchmark for NLP knowledge mining, with its full-text annotation paradigm and fine-grained ontology design enabling downstream applications such as domain-specific QA and technological trend analysis.

### v4

@@ Abstract
Structured information extraction from scientific literature is crucial for capturing core concepts and emerging trends in specialized fields. While existing datasets aid model development, most focus on specific publication sections due to domain complexity and the high cost of annotating scientific texts. To address this limitation, we introduce SciNLP—a specialized benchmark for full-text entity and relation extraction in the Natural Language Processing (NLP) domain. The dataset comprises 60 manually annotated full-text NLP publications, covering 6,409 entities and 1,648 relations. Compared to existing research, SciNLP is the first dataset providing full-text annotations of entities and their relationships in the NLP domain. To validate the effectiveness of SciNLP, we conducted comparative experiments with similar datasets and evaluated the performance of state-of-the-art supervised models on this dataset. Results reveal varying extraction capabilities of existing models across academic texts of different lengths. Cross-comparisons with existing datasets show that SciNLP achieves significant performance improvements on certain baseline models. Using models trained on SciNLP, we implemented automatic construction of a fine-grained knowledge graph for the NLP domain. Our KG has an average node degree of 3.3 per entity, indicating rich semantic topological information that enhances downstream applications. The dataset is publicly available at: https://github.com/AKADDC/SciNLP.
@@ 1 Introduction
Scientific Named Entity Recognition (SciNER) and Scientific Relation Extraction (SciRE) have emerged as two pivotal technologies for scientific literature mining. SciNER focuses on identifying and categorizing domain-specific scientific entities in publications(Jeong and Kim, 2022; Zhang et al., 2021; Zhong and Chen, 2021b; Beltagy et al., 2019), while SciRE detects semantic relationships between entities and classifies them into predefined relation types(Yan et al., 2021; Chen et al., 2020; Ji et al., 2020; Wadden et al., 2019). These tasks play essential roles in downstream applications such as scientific leaderboard construction, academic Q&A systems, knowledge graph building, text summarization, and method recommendation(Wang et al., 2024; Cao et al., 2021; Zhang et al., 2024a; Kedia et al., 2021).
@@
We construct SciNLP, the first manually curated, full-text scientific publication dataset for the NLP domain, comprising 60 ACL conference long papers (2001–2024) with comprehensive semantic annotations. The dataset includes 6,409 fine-grained entities and 1,648 complex relations, addressing the limitations of existing scientific information extraction resources confined to abstracts or narrow subdomains. SciNLP pioneers systematic full-text entity annotation for NLP literature.
@@ 2.2 Scientific Entity Relation Extraction
Reviewing the extant annotated datasets reveals significant imbalances in disciplinary coverage and annotation depth across current scientific publication resources for SciIE. Cross-domain corpora such as predominantly concentrate on AI and ML domains(Luan et al., 2018; Pan et al., 2024; Zhang et al., 2024b), while specialized disciplinary datasets exhibit marked fragmentation in textual coverage, typically constrained to specific paper sections or isolated sentences(Jain et al., 2020; Gábor et al., 2018), thereby compromising the integrity of holistic scientific discourse. Addressing these limitations, this study proposes the creation of SciNLP – the inaugural full-text annotation benchmark for NLP research. Anchored in comprehensive semantic units from 60 ACL proceedings (2001-2024), the corpus achieves granular annotation of 6,409 entities and 1,648 relations. Furthermore, it establishes a sophisticated taxonomy of entity and relational categories specifically tailored to the structural and conceptual composition of NLP scholarly articles.
@@ 3.2 Comparison with Previous Datasets
Table 2: Examples of the relations types, the four entity types: Model, Metric, Dataset, Task.
@@ 3.3 Data Analysis
Upon completing the annotation process, our dataset comprises 6,409 fine-grained entities and 1,648 relations, with an average of 106 entities and 27 relations per document. As shown in Table 1, the scale of our dataset substantially exceeds prior NLP-specific datasets designed for entity and relation extraction. The annotated data was then randomly shuffled and split into training, validation, and testing sets in an 8:1:1 ratio to support subsequent experimental analyses.
@@ 4.2 Evaluation Settings and Results
Boundary-Aware Metric (Rel): Following prior work(Zhong and Chen, 2021a; Ye et al., 2022; Yan et al., 2023), a relation is considered correct only if: (a) Subject entity boundaries are correctly identified; (b) Object entity boundaries are correctly identified; (c) Relation type is accurately predicted.
@@
PURE(Zhong and Chen, 2021a) 59.42 57.51 56.15 57.03 60.25 59.93 58.35 59.93
@@
HGERE 79.53 49.28 47.64 - 74.91 55.72 43.6 - 86.85 62.32 61.10 -
Table 4: F1 scores of different baseline models on the SciNLP, SciERC, and SciER datasets.
@@ 5.1 Supervised Baselines
PURE: (Zhong and Chen, 2021a) This model employs a dual-encoder architecture to decouple entity recognition and relation extraction tasks. The entity model performs span-based type prediction, while the relation model enhances classification through dynamic entity pair inputs and the insertion of entity-type markers. It introduces crosssentence context expansion by extending input windows, leveraging pre-trained models’ ability to capture long-range dependencies for improved global semantic reasoning.
@@ 5.2 Result Analysis
Table  4 presents a comparative evaluation of different baseline models on three scientific information extraction datasets: SciNLP, SCIERC, and SciER, revealing notable interaction effects between dataset characteristics and model architectures.
First, for the named entity recognition (NER) task, all three datasets exhibit a consistent performance improvement as model complexity increases. Taking HGERE as an example, it significantly outperforms PURE and PL-Marker across SciNLP, SCIERC, and SciER, indicating that the incorporation of graph-structured modeling and global semantic interaction mechanisms effectively enhances the representation of complex scientific entities. Meanwhile, the overall NER performance on SciNLP remains lower than that on SciER but is comparable to SCIERC, suggesting that entity recognition performance is still largely influenced by annotation scale and domain consistency.
Second, more pronounced performance disparities emerge in relation extraction–related tasks. Overall, SciER consistently achieves superior performance on both the Rel and Rel+ tasks compared to SciNLP and SCIERC, which can be attributed to its higher relation annotation density and more mature dataset construction paradigm. Notably, on the SciNLP dataset, HGERE performs substantially worse on the Rel and Rel+ tasks than on SCIERC and SciER, contradicting earlier assumptions that it would demonstrate a clear advantage on SciNLP. This observation indicates that HGERE’s graph-based reasoning capability is not fully exploited in SciNLP’s relation modeling setting.
In summary, while SciNLP effectively supports the training of complex models for entity recognition, its current annotation density limits the extent to which graph-based models can fully realize their potential advantages in relation modeling, pointing to a clear direction for future improvements in dataset design.
@@ 6 Building Fine-grained KG in NLP Domain
We construct a scientific knowledge graph from the entire ACL corpus. First, we built a comprehensive corpus containing full texts and metadata of all ACL conference papers (2001–2024) from the ACL Anthology repository, totaling 82,672 papers. Leveraging the HGERE and PL-Marker models, we performed end-to-end triplet extraction across all full-text papers to build a fine-grained NLP knowledge graph, where nodes represent extracted scientific entities and edges denote semantic relationships between entities. This framework enables systematic encoding of domain-specific knowledge structures and cross-document scientific logic.
The construction of our knowledge graph involves two key steps: First, the model extracts all triplet data T from each paper, preserving every triplet in the form ($e_{i}$, $r_{k}$, $e_{j}$), where $e_{i}$ and $e_{j}$ are related entities, and $r_{k}$ denotes their semantic relationship. Second, we perform entity normalization to ensure uniqueness within the knowledge graph. This process begins by creating an abbreviation-full form mapping table between entities using the predefined similar relationship. Subsequently, we unify entities using the entity normalization method proposed by(Zhang et al., 2024a), which calculates similarity scores based on lexical features and applies clustering algorithms to standardize entity representations. This dual-step approach guarantees both the richness and semantic consistency of the knowledge graph. In the end, we constructed a fine-grained NLP knowledge graph comprising 205k nodes and 693k edges, with an average node degree of 3.3. Figure  2 illustrates a subset of the constructed graph, demonstrating its ability to effectively preserve domain-specific knowledge in NLP. This topology reflects the intricate semantic networks inherent to computational linguistics research.
@@
Task 33,684 measuredBy 117,050
Dataset 31,587 evaluatedBy 24,325
Metric 30,751 usedFor 171,930
Model 109,636 enhancedBy 32,329
evaluatedOn 37,742
partOf 44,462
compareWith 77,039
subtaskOf 40,454
similarWith 37,951
trainedWith 91,147
subclassOf 18,822
Total 205,658 Total 693,251
@@ 7 Conclusion
We present SciNLP, a dataset specifically designed for fine-grained entity and relation extraction in the Natural Language Processing (NLP) domain. The dataset defines four entity types (Task, Method, Dataset, Metric) and 11 relation types (MeasuredBy, UsedFor, TrainedWith, EnhancedBy, SubclassOf, SubtaskOf, PartOf, EvaluatedBy, EvaluatedOn, SimilarWith, CompareWith), overcoming the limitations of coarse-grained relation definitions in existing resources. Constructed from 60 ACL conference papers spanning 2001–2024, SciNLP contains 6,409 entities and 1,648 relations. Unlike cross-disciplinary or abstract-level datasets, SciNLP preserves full-text scientific context through document-level annotation, providing critical support for models to capture cross-paragraph semantic dependencies. Experimental results demonstrate that models trained on SciNLP significantly outperform sentence-level settings in document-context modeling. For instance, the PL-Marker model achieves a 17.62 point improvement on the strict relation classification (Rel+) task compared to the cross-disciplinary SciERC dataset, validating the enhanced capability of domain-specific annotation for complex relation reasoning. Leveraging SciNLP, we automatically extracted knowledge from 82,672 NLP publications, constructing a fine-grained knowledge graph with 205k nodes and 693k edges, where entities exhibit an average node degree of 3.3. SciNLP establishes a high-confidence benchmark for NLP knowledge mining, with its full-text annotation paradigm and fine-grained ontology design enabling downstream applications such as domain-specific QA and technological trend analysis.